A rental car costs a minimum charge plus a fixed charge per mile driven. One customer was charged $60 for 160 miles of use. Another customer drove 280 miles and paid $90. Write an equation relating the total cost, C (in dollars), to the miles driven, x.
Answer by partha_ban(41)

Assume m = minimum charge
and r = fixed charge per mile
Therefore, C = m + r.x ... ... ...(A)
where C = Total cost
and x = miles driven
So, 60 = m + 160r ... ... ...(1)
and, 90 = m + 280r ... ... ...(2)
Subtracting (1) from (2),
120r = 30
r = 30/120 = 1/4
Substituting r = 1/4 in (1),
m + 160*1/4 = 60
m + 40 = 60
m = 60 - 40
m = 20
Therefore from (A):
C = 20 + x*1/4
C = 20 + x/4